can you help me solve? what is the solution set?
t(5t+2)=0 (need solution set)
Thank you!




Answer by nerdybill(7384)

t(5t+2)=0
there are two terms:
t
and
5t+2
.
Set EACH to zero to give you the solution set:
t = 0 (first solution)
.
5t+2 = 0
5t = -2
t = -2/5 (second solution)
.
Solution set is then:
t = {-2/5, 0}